Circular No. 7287

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7244/7440/7444 (for emergency use only)

COMET C/1999 U1 (FERRIS) B. Koehn, Lowell Observatory, adds that the discovery images (cf.

IAUC 7283) show this object as clearly cometary, with a coma diameter of ∼ 9′′ and a fan-shaped tail extending 18′′ from the comet toward the south. Additional astrometry and the following preliminary parabolic orbital elements appear on MPEC 1999-U19:

T = 1998 Aug. 17.978 TT  ω = 268.^{o}285
Ω = 59.719 }2000.0
q = 2.92087 AU  i = 96.433
1999TT α_{2000}  δ_{2000}  Δ  r  ϵ  β  m_{1}
Oct. 19 4^{h}38.^{m}14 + 8^{o}26.^{′}8 4.242 4.990 134.^{o}4 8.^{o}2 16.6
29 4 29.08 + 8 56.4 4.206 5.060 146.0 6.3 16.7
Nov. 8 4 19.00 + 9 27.8 4.202 5.131 157.4 4.3 16.7
18 4 08.36 +10 01.0 4.232 5.201 167.3 2.4 16.8
28 3 57.69 +10 36.1 4.301 5.272 168.8 2.1 16.9
Dec. 8 3 47.50 +11 12.8 4.407 5.343 159.9 3.6 17.0

NOVA IN IC 1613 J. Y. King, M. Modjaz, and W. D. Li, University of California at

Berkeley, on behalf of the Lick Observatory Supernova Search, report the discovery with the 0.8-m Katzman Automatic Imaging Telescope (KAIT) of an apparent nova in IC 1613. The nova was discovered on unfiltered images taken on Oct. 12.3 UT (mag ∼ 18.6), and confirmed on unfiltered images taken on Oct. 13.3 (mag ∼ 18.8). Inspection of the KAIT archive data shows that the nova has been in KAIT images since Aug. 23, all with magnitudes ∼ 18.5. The new object is located at α = 1^{h}04^{m}48^{s}.97, δ = +2^{o}05′28′′.8 (equinox 2000.0). A KAIT image of the same field on Aug. 18 (limiting mag ∼ 19.0) showed nothing at the position of the nova. IC 1613 is a dwarf galaxy in the Local Group.

NOVAE IN THE SMALL MAGELLANIC CLOUD Clarifying the title on IAUC 7286, it is to be noted that the novae

reported on IAUC 7121 and 7239 should be called ‘Nova in the Small Magellanic Cloud 1999 No. 1’ and ‘Nova in the Small Magellanic Cloud 1999 No. 2’, respectively.

1999 October 21  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1999 CBAT  Daniel W. E. Green